OpenStudy (anonymous):

Math test... Three guys go to get a room for the night. The room is 30 bux. Each guy gives ten bux. Get the key and turn in for the night. Shortly after, the clerk realizes the room is only 25 dollars. He sends a bellboy to the room with five dollars. He can't figure out how to divide 5 bux amongst the three men, so he gives each guy a dollar and keeps two bux as his tip. So....each guy paid ten dollars, but got one dollar back...so each guy paid 9 dollars. The room was originally 30 dollars. 9 times 3 is 27. Two dollar tip. 27+2=29..... Where's the other dollar?

OpenStudy (amistre64):

you might want to try solving it correctly instead of making up bad mathing solutions for it

OpenStudy (amistre64):

each guy paid 9 dollars, for a total of 27 dollars for a 25 dollar room, 25-27 = 2 .... where did the 2 dollars go?

OpenStudy (anonymous):

Tip

OpenStudy (amistre64):

then there is no paradox, or conundrum, or mystery about it when you do the math correctly.

OpenStudy (amistre64):

(-10-10-10) + 30 = 0 (-10+1-10+1-10+1) + 30-5 = 0+1+1+1-5 (-9-9-9) + 25 = -2 -27 + 25 = -2

OpenStudy (amistre64):

if we run the scenario in the post: (-10-10-10) + 30 = 0 (-10+1-10+1-10+1) -2 +30 = 0 (-9-9-9) -2 +30 = 0 -29 + 30 = 0 its just not math .... its make believe :)

OpenStudy (anonymous):

Ok thanks